The following information was derived from an Income Statement when 20,000 units were sold.

Sales $1,250,000
Variable Costs 250,000
Contribution Margin $1,000,000
Fixed Costs 400,000
Net Income $ 600,000

Determine the break-even point in units sold.

What is the margin of safety?

How many units must be sold if a Net Income of $800,000 is desired?
Answer #1

Contribution Margin per unit=$1,000,000/20,000

=$50 per unit

a.Hence breakeven=Fixed Cost/Contribution Margin

=(400,000/50)

=8,000 units.

b.Margin of safety=Total sales-Breakeven sales

=(20,000-8000)

=12,000 units.

c.Target contribution margin=Fixed cost+Target profits

=(400,000+800,000)=$1,200,000

Hence units to be sold=$1,200,000/50

=24,000 units.